OpenStudy (anonymous):

Write the slope-intercept form of the equation parallel to y = 7x + 2, which passes through the point (1, -3). A. y = 7x - 10 B. y = -1/7 x + 2 C. y = 7x - 3 D. y = -7x + 10

OpenStudy (jdoe0001):

2 parallel lines have THE SAME SLOPE

OpenStudy (jdoe0001):

so by looking at the y = mx+b point-slope form from the one above we get that the slope is \(\large y = \color{red}{7}x+2\)

OpenStudy (jdoe0001):

now to get the slope-intercept form, which is just the expanded version of that, just do http://www.teachhandson.com/calculationb.gif

OpenStudy (jdoe0001):

you have a point (1, -3), you have the slope, plug them in, solve for "y" :)

OpenStudy (anonymous):

Im so confused

OpenStudy (jdoe0001):

ok, since both lines are parallel, the GIVEN one, and the one with the point (1, -3), so both have the same slope. parallel lines have the same slopes

OpenStudy (anonymous):

Okay

OpenStudy (jdoe0001):

so by looking at the y = mx+b point-slope form from the one above we get that the slope is \(\large y = \color{red}{7}x+2\)

OpenStudy (anonymous):

Okay.. But there is no equation above for the answer

OpenStudy (jdoe0001):

so, the slope for the other line, will also be "7", now it has a point (1, -3) (1, -3) (x, y) now plugging those 2 in the point-slope form, we get $$ (x,y) \implies (\color{red}{1},\color{blue}{-3})\\ y-(\color{blue}{-3})=7(x-\color{red}{1}) $$

OpenStudy (jdoe0001):

now you just need to solve for "y" to get the slope-point intercept form

OpenStudy (anonymous):

Would it be B then?

OpenStudy (jdoe0001):

well, what did you get for "y"?